OpenStudy (anonymous):

Show that the equation \[\frac{x^5}{120}+\frac{x^4}{24}+\frac{x^3}{6}+\frac{x^2}{2}+x+1=0\]has no repeated roots.

OpenStudy (shubhamsrg):

the ques is quite similar to Q.9 of http://www.cmi.ac.in/admissions/sample-qp/ugmath2010.pdf i have the solution and that might sound interesting,,i just solved this paper yesterday! :D

OpenStudy (kainui):

Looks like it has something to do with its derivatives.

OpenStudy (shubhamsrg):

if f(x) = x^5/5! + x^4/4! ...... x +1 now x=0 is not a solution for f(x)=0.. then f'(x) = x^4/4! + x^3/3! .... x + 1 i.e. f'(x) = f(x) - x^5/5! .................(1) now let f(x) = ( (x-a)^m ) * g(x) where m is some natural no. > 1 it means a is a repeated root of f(x) now f'(x)= ( m(x-a)^(m-1) ) *g(x) + g'(x) * ( (x-a)^m ) = (x-a) * (something) hence a is also a root of f'(x) coming back to eqn (1), for x= a , we have f'(a) = f(a) - a^5/5! =>0 = 0 - a^5 /5! =>a=0 which states a=0 is the repeated root but its a contradiction as x=0 is not a solution for f(x) =0 hence no such a exists or in other words,,f(x) has no repeated roots!!

OpenStudy (shubhamsrg):

for soln of q.9 of that pdf file,,you just generalize ..

OpenStudy (anonymous):

nice work... @shubhamsrg let \(x=a\) be possible real root of equation \(f(x)=0\) we know that \(a\neq 0\) and\[f'(a)=-\frac{a^5}{120} \neq 0\]